OpenStudy (gigi1103):

Which related equation can be used to solve the equation? 1. 8 + a = 3.9 A. a = –3.9 – 8 2. 3.9 + a = 8 B. a = 3.9 – 8 3. 3.9 + a = –8 C. a = 8 – 3.9 4. 8 + a = –3.9

OpenStudy (gigi1103):

@Nnesha

OpenStudy (gigi1103):

Which related equation can be used to solve the equation? Column A Column B 1. 8 + a = 3.9 2. 3.9 + a = 8 3. 3.9 + a = –8 4. 8 + a = –3.9 A. a = –3.9 – 8 B. a = 3.9 – 8 C. a = 8 – 3.9